# Robust Finetuning of Vision-Language-Action Robot Policies via Parameter Merging

### v2

@@ 4 Challenges of Finetuning Generalist Robot Policies
To understand the challenges of robust finetuning of generalist robot policies, we start by evaluating the standard finetuning approach. We evaluate finetuning the full model (Task-FT) in LIBERO [31], a multi-task robotic manipulation simulator containing 130 total tasks. Concretely, given a state-of-the-art vision-language-action policy [8], pretrained on demonstration data from 117 tasks from the LIBERO-{90, goal, spatial, object} suites, we finetune it to a new LIBERO target task. We consider three target tasks from LIBERO-10: mugs-on-plates, pot-on-stove, and items-into-basket. We then measure performance in the three scenarios introduced in Section˜3: ID, OOD, and Generalist. For OOD evaluations, we alter object positions, add new distractors, and change backgrounds. Generalist evaluations are performed over $20$ tasks from the pretraining dataset. More details about the OOD evaluations in Section˜6.1 and Section˜A.6.
@@
Fig.˜4 shows the performance of Task-FT on three types of evaluations. We find a clear tradeoff between generalist and ID target task performance: though the model gets better in ID target task after fine-tuning, it increasingly loses its generalist capabilities. Additionally, when the model is finetuned for too long, it even starts losing performance for ID tasks. Both of these phenomena are likely because the model has severely overfitted to the small demonstration dataset, and is unable to do other tasks or recover from small mistakes unseen in the dataset. More importantly, one would hope that the finetuned model can generalize to small variations of the target task not seen in the finetuning dataset, since the pretrained model contains knowledge for such generalization. However, all tested checkpoints above show a large gap between ID and OOD performance, showing that the model is not able to complete the target task when there are small variations present. Section˜A.3 shows that even with careful tuning of learning rate and number of gradient steps, such overfitting still exists.
@@ 6.1 Experimental Setup
We evaluate RETAIN in real-world and simulated finetuning settings (see Fig.˜6). For our real-world experiments, we use the DROID robot setup [25], which consists of a 7-DoF Franka robot arm with a wrist-mounted camera and at least one external camera. We design two challenging fine-tuning tasks: wiping the whiteboard with an eraser (which we call whiteboard) and putting the dishes into a drying rack (which we call plates). For both tasks, we collect roughly $50$ and $100$ demonstrations, respectively. All demonstrations are collected in a single environment with minor position variations (see Fig.˜6, blue), to mirror the variation in typical narrow-data finetuning regimes [71]. We test generalization of the finetuned policies on a much broader set of target environments (Fig.˜6, yellow), which include unseen backgrounds, object instances, and camera views. See example task trajectories in Fig.˜3. For our simulated experiments, we use the LIBERO simulation environment [31]. We finetune policies pretrained on the LIBERO-{object, spatial, goal, 90} datasets to three new tasks from the LIBERO-10 suite: pot-on-stove, mugs-on-plates, and items-into-basket. We use $\approx 45$ demonstrations per task, obtained after filtering and preprocessing the 50 demos provided with the LIBERO simulator, which only contain minor variations to the initial positions of each object, and again test on a much broader distribution of initial positions, backgrounds, and additional distractors (see Fig.˜6, right). More setup details can be found in Section˜A.5 and Section˜A.6.
@@
We compare against prior methods, including those that incorporate regularization techniques to reduce overfitting, in the robust finetuning setting. Specifically, we compare our approach, RETAIN, to: “Task-FT”, which finetunes the pretrained policy only on the target task dataset using behavioral cloning (Eq.˜1) [6, 8]; “Co-FT”, which finetunes on a mix of pretraining and target task data to reduce overfitting [15, 12], “LoRA”, which uses low-rank adaptation [17] during finetuning to retain more of the pretraining capabilities [41, 26]; “Freeze-FT”, which freezes the language model backbone during finetuning and only updates the vision encoder, and, in the case of $\pi_{0}$, the action expert output head, following similar approaches in prior work, e.g., Kim et al. [26], Zhang et al. [71]; Scratch, which learns a policy from scratch on the demonstration dataset instead of finetuning a generalist policy. For more details about the policy classes and implementations, see Section˜A.8. Section˜A.7 details our choice of hyperparameters and tuning process.
@@ 6.3 RETAIN Scales with the Amount of Pretraining Data
As we saw in Section˜6.2, RETAIN is not only able to perform well on tasks from the pretraining distribution, but also solve broader variations of finetuning tasks, as if it “merged” the generalist ability of the pretrained model with the task specialization of the finetuned model. This begs the question: how well does RETAIN do when the pretrained model contains differing amounts of “generalist knowledge”, as induced by different amounts of pretraining data? To study how differing amounts of pretraining data affects RETAIN’s performance, we use the DROID dataset, containing a large quantity of diverse robot data. We consider three pretrained generalist VLA policies: (1) the public $\pi_{0}$-FAST-DROID checkpoint that is trained on all of DROID and a large repertoire of robot data from Physical Intelligence [45], (2) only trained on all of DROID, totaling $76k$ episodes, and (3) only trained on the subset of DROID collected at Berkeley and Stanford, totaling $20k$ episodes. All three policies are pretrained by taking roughly $1$ epoch over the dataset, and finetuned on the plates task with the same hyperparameters. We then evaluate the three RETAIN-co-FT policies, each obtained by merging the respective pretrained policy with its finetuned counterpart, on ID and OOD scenes in Fig.˜9. In ID evaluations, DROID-All + PI performs the best, slightly outperforming the two other models pretrained with less data. In OOD evaluations, more pretraining data leads to significantly better performance, with the best model performing nearly as good on OOD evaluation as it does on ID evaluation. This shows that RETAIN scales with the amount of pretraining data, and can better “transfer” generalist knowledge from the pretrained model to new scenes when the pretrained model is more general.
@@ 6.4 Analyzing the Importance of Merging Different Parameters in RETAIN
Figure 11: Language model parameters have the most influence in modality-specific merging. Left: Merged model’s OOD performance over a grid search of $\alpha_{a},\alpha_{v},\alpha_{l}$, and $\alpha_{l}$ has the most impact. Middle: OOD performance of $\alpha_{a}$ and $\alpha_{v}$ averaged over different $\alpha_{l}$, and higher values are better. Right: Merging only the language model parameters ($\alpha_{a}=\alpha_{v}=1,\alpha_{l}<1$) improves performance over a uniform coefficient $\alpha$ for all parameters.
@@ 8 Limitations
While we empirically verified that RETAIN works exceptionally well in helping the finetuned model generalize to out-of-distribution variations of the task, we don’t understand the full scope of the reasons why model parameter merging was able to lead to such generalization. This is an interesting area for future work. We have included some discussion in Section˜A.9 of some hypothesess and why previous work found model parameter merging effective for vision and language tasks. Additionally, RETAIN involves an important hyperparameter, the merginge coefficient, that can be tuned. While we find in our real world experiments that RETAIN is robust to different values of this parameter, slight tuning of this parameter is needed. One avenue of future work is determining a good heuristic of how to choose this value.
@@ A.2 LIBERO Results Details
A.3 Ablation on Learning Rate and Gradient Steps
Here, we ablate the learning rate and number of gradient steps we take in the task-FT policy in Fig.˜4 to study whether better hyperparameter choices can reduce or resolve overfitting [^1] Following the best practice from Black et al. [8], we always use a learning rate warmup period of $1000$ steps in LIBERO and a consine decay schedule over $30k$ steps to $1/10$ of the peak learning rate. The learning rate we report here is the peak learning rate. . In Fig.˜4, we use learning rate 2.5e52.5e-5. Here in Fig.˜15, we ablate four different learning rates: one greater than the original and two smaller. We evaluate the model at every $100$ gradient steps to also ablate on the number of gradient steps we take. With a larger learning rate, it’s clear that the overfitting issue is more severe, and the performance on all three kinds of evaluations (ID, OOD, Generalist) go down to near zero. With a smaller learning rate, the model still performs well in ID evaluations, and suffers less from forgetting generalist capabilities (measured from Generalist evaluations). This is to be expected because the finetuned model is closer to the pretrained model with a smaller learning rate. However, note that with a smaller learning rate, the OOD evaluation performance is also worse. We plot the OOD performance achieved using the original learning rate in Fig.˜4 as the dotted orange line in Fig.˜15, and the gap between the dotted and solid orange line shows the performance gap in OOD evaluation when we lower the learning rate, possibly due to underfitting. This experiment shows that tuning the learning rate and the number of update steps does not solve the problem of overfitting during finetuning. While lowering the learning rate can retain more generalist knowledge, it does not prevent the gradual loss of it. More importantly, the lower learning rate leads to worse OOD evaluation performance.
@@ A.3 Ablation on Learning Rate and Gradient Steps
A.4 Analysis of Finetuning Path in Parameter Space
@@
To start out, we check the colinearity of the vectors $\theta_{i+1}-\theta_{i}$ and $\theta_{i}-\theta_{i-1}$, where $\theta_{i}$ is the parameter of the finetuned checkpoint at gradient steps 100i100*i (i.e. we plot the difference vector every $100$ steps during finetuning). We measure the colinearity as the cosine similarity between the two difference vectors. Fig.˜16 shows this for the four different learning rates we ablated in Fig.˜15. Since no values are close to $1$, this shows that the changes in parameter space is highly non-linear. This is expected since the parameter trajectory of deep neural networks is often highly non-linear.
@@
Next, we attempt to more directly visualize the path/direction of the parameters during finetuning by projecting them down to 2D with Principal Component Analysis (PCA). Specifically, we again consider the difference of the weight vectors, $X_{i}=\theta_{i+1}-\theta_{i}$, at every $100$ steps during finetuning. In Fig.˜17, we plot the first two principal components of $X_{i}$ in blue, and label the points $i$. Indeed, we see that for all learning rates, the direction of the parameters is highly non-linear. With small learning rates, the direction oscillates a lot; with larger learning rates, the direction bends in a certain direction. In addition, we plot the two principal components of the parameter-merged model as well in orange. As expected, since the the merged model takes a linear path and the finetuned one does not, the two models end in in very different places in the parameter space. This shows that model merging achieves a different solution than any checkpoints on the finetuning path.
Finally, we analyze changes in all directions of the difference vectors $X_{i}$, instead of just the two principal components as shown in Fig.˜17. We take the difference vector matrix $Y=[X_{1};X_{2};...,X_{n}]$ and compute the singular values of $YY^{T}$ and Fig.˜18. If the parameter vectors lie in a linear path, then all difference vectors would point in the same direction and there should only be $1$ non-zero singular value. However, it’s clear from Fig.˜18 that most singular values are non-zero, showing that the path is non-linear in many dimensions. This generalizes the intuition from Fig.˜17 to more dimensions, and shows that model merging indeed achieves a different solution than the finetuning path.
@@
A.5 DROID Setup Details
@@ A.5.4 Evaluation Details
A.5.5 Evaluation Criteria: Whiteboard
@@
A.5.6 Evaluation Criteria: Plates
@@
A.5.7 Generalist Evaluations Details
@@
A.6 LIBERO Setup Details
@@
A.6.1 LIBERO Pretraining
In order to obtain a base-model to serve as the starting point for RETAIN in LIBERO, we pretrain $\pi_{0}$ on a mixture of LIBERO datasets. Specifically, we use 90 tasks from LIBERO-90, 9 tasks from LIBERO-object, 9 tasks from LIBERO-spatial, and 9 tasks from LIBERO-goal, for a total of 117 tasks in our pretraining dataset. For all LIBERO datasets, pretraining and finetuning, we utilize pre-processed RLDS versions gathered from here and here. These datasets consist of LIBERO demonstrations that have been preprocessed to upscale images, filter out transitions with idle actions, and remove failure trajectories. More details, such as hyperparameter choices, for the pretraining stage itself can be found in A.7. All subsequent fine-tuning and evaluation uses the normalization stats of this pretraining dataset, applying mean/standard-deviation normalization.
A.6.2 LIBERO Finetuning Datasets
As mentioned in 6.1, we finetune on 3 tasks from LIBERO-10: pot-on-stove, mugs-on-plates, and items-into-basket. For each dataset, we obtain pre-processed and filtered versions as described above. 19 highlights what these finetuning tasks, and thus also our ID evals, look like.

### v3

# arXiv:2512.08333v3[cs.RO] 27 Feb 2026

@@ 4 Challenges of Finetuning Generalist Robot Policies
To understand the challenges of robust finetuning of generalist robot policies, we start by evaluating the standard finetuning approach. We evaluate finetuning the full model (Task-FT) in LIBERO [31], a multi-task robotic manipulation simulator containing 130 total tasks. Concretely, given a state-of-the-art vision-language-action policy [8], pretrained on demonstration data from 117 tasks from the LIBERO-{90, goal, spatial, object} suites, we finetune it to a new LIBERO target task. We consider three target tasks from LIBERO-10: mugs-on-plates, pot-on-stove, and items-into-basket. We then measure performance in the three scenarios introduced in Section˜3: ID, OOD, and Generalist. For OOD evaluations, we alter object positions, add new distractors, and change backgrounds. Generalist evaluations are performed over $20$ tasks from the pretraining dataset. More details about the OOD evaluations in Section˜6.1 and Section˜A.7.
@@
Fig.˜4 shows the performance of Task-FT on three types of evaluations. We find a clear tradeoff between generalist and ID target task performance: though the model gets better in ID target task after fine-tuning, it increasingly loses its generalist capabilities. Additionally, when the model is finetuned for too long, it even starts losing performance for ID tasks. Both of these phenomena are likely because the model has severely overfitted to the small demonstration dataset, and is unable to do other tasks or recover from small mistakes unseen in the dataset. More importantly, one would hope that the finetuned model can generalize to small variations of the target task not seen in the finetuning dataset, since the pretrained model contains knowledge for such generalization. However, all tested checkpoints above show a large gap between ID and OOD performance, showing that the model is not able to complete the target task when there are small variations present. Section˜A.4 shows that even with careful tuning of learning rate and number of gradient steps, such overfitting still exists.
@@ 6.1 Experimental Setup
We evaluate RETAIN in real-world and simulated finetuning settings (see Fig.˜6). For our real-world experiments, we use the DROID robot setup [25], which consists of a 7-DoF Franka robot arm with a wrist-mounted camera and at least one external camera. We design two challenging fine-tuning tasks: wiping the whiteboard with an eraser (which we call whiteboard) and putting the dishes into a drying rack (which we call plates). For both tasks, we collect roughly $50$ and $100$ demonstrations, respectively. All demonstrations are collected in a single environment with minor position variations (see Fig.˜6, blue), to mirror the variation in typical narrow-data finetuning regimes [71]. We test generalization of the finetuned policies on a much broader set of target environments (Fig.˜6, yellow), which include unseen backgrounds, object instances, and camera views. See example task trajectories in Fig.˜3. For our simulated experiments, we use the LIBERO simulation environment [31]. We finetune policies pretrained on the LIBERO-{object, spatial, goal, 90} datasets to three new tasks from the LIBERO-10 suite: pot-on-stove, mugs-on-plates, and items-into-basket. We use $\approx 45$ demonstrations per task, obtained after filtering and preprocessing the 50 demos provided with the LIBERO simulator, which only contain minor variations to the initial positions of each object, and again test on a much broader distribution of initial positions, backgrounds, and additional distractors (see Fig.˜6, right). More setup details can be found in Section˜A.6 and Section˜A.7.
@@
We compare against prior methods, including those that incorporate regularization techniques to reduce overfitting, in the robust finetuning setting. Specifically, we compare our approach, RETAIN, to: “Task-FT”, which finetunes the pretrained policy only on the target task dataset using behavioral cloning (Eq.˜1) [6, 8]; “Co-FT”, which finetunes on a mix of pretraining and target task data to reduce overfitting [15, 12], “LoRA”, which uses low-rank adaptation [17] during finetuning to retain more of the pretraining capabilities [41, 26]; “Freeze-FT”, which freezes the language model backbone during finetuning and only updates the vision encoder, and, in the case of $\pi_{0}$, the action expert output head, following similar approaches in prior work, e.g., Kim et al. [26], Zhang et al. [71]; Scratch, which learns a policy from scratch on the demonstration dataset instead of finetuning a generalist policy. For more details about the policy classes and implementations, see Section˜A.9. Section˜A.8 details our choice of hyperparameters and tuning process.
@@ 6.3 RETAIN Scales with the Amount of Pretraining Data
As we saw in Section˜6.2, RETAIN is not only able to perform well on tasks from the pretraining distribution, but also solve broader variations of finetuning tasks, as if it “merged” the generalist ability of the pretrained model with the task specialization of the finetuned model. This begs the question: how well does RETAIN do when the pretrained model contains differing amounts of “generalist knowledge”, as induced by different amounts of pretraining data? To study how differing amounts of pretraining data affects RETAIN’s performance, we use the DROID dataset, containing a large quantity of diverse robot data. We consider three pretrained generalist VLA policies: (1) the public $\pi_{0}$-FAST-DROID checkpoint that is trained on all of DROID and a large repertoire of robot data from Physical Intelligence [45], (2) only trained on all of DROID, totaling $76k$ episodes, and (3) only trained on the subset of DROID collected at Berkeley and Stanford, totaling $20k$ episodes. All three policies are pretrained by taking roughly $1$ epoch over the dataset, and finetuned on the plates task with the same hyperparameters. We then evaluate the three RETAIN-co-FT policies, each obtained by merging the respective pretrained policy with its finetuned counterpart, on ID and OOD scenes in Fig.˜9. To account for the performance difference of the base policies, we report the success rate increase between RETAIN-co-FT and co-FT. In ID evaluations, policies trained with differing amount of data perform similarly. In OOD evaluations, more pretraining data leads to significantly better performance. This shows that RETAIN scales with the amount of pretraining data, and can better “transfer” generalist knowledge from the pretrained model to new scenes when the pretrained model is more general. In fact, the best performing model (DROID-ALL + PI) performs nearly as good on OOD evlauation as it does on ID evaluation, as shown in Fig.˜15.
@@ 6.4 Analyzing the Importance of Merging Different Parameters in RETAIN
Figure 11: Language model parameters have the most influence in modality-specific merging. Left: Merged model’s OOD performance over a grid search of $\alpha_{a},\alpha_{v},\alpha_{l}$, and $\alpha_{l}$ has the most impact. Middle: OOD performance of $\alpha_{a}$ and $\alpha_{v}$ averaged over different $\alpha_{l}$, and higher values are better. Right: Merging only the language model parameters ($\alpha_{a}=\alpha_{v}=1,\alpha_{l}<1$) performs similarly to merging all parameters.
@@ 8 Limitations
While we empirically verified that RETAIN works exceptionally well in helping the finetuned model generalize to out-of-distribution variations of the task, we don’t understand the full scope of the reasons why model parameter merging was able to lead to such generalization. This is an interesting area for future work. We have included some discussion in Section˜A.10 of some hypothesess and why previous work found model parameter merging effective for vision and language tasks. Additionally, RETAIN involves an important hyperparameter, the merginge coefficient, that can be tuned. While we find in our real world experiments that RETAIN is robust to different values of this parameter, slight tuning of this parameter is needed. One avenue of future work is determining a good heuristic of how to choose this value.
9 Acknowledgements
This research was partly supported by ONR N00014-25-1-2060. This research used the Savio computational cluster resource provided by the Berkeley Research Computing program at UC Berkeley. We would also like to thank the anonymous ICLR reviewers for thoughtful comments and suggestions.
@@ A.2 LIBERO Results Details
A.3 Data Scaling Results Details
Fig.˜15 highlights more details of the data scaling results, examining for which base models RETAIN provides the greatest benefit: it shows the ID/OOD performance for co-FT and RETAIN-co-FT with varying amounts of pretraining data that led to the difference results presented in Fig.˜9. We see that as we scale pretraining data, ID performance is similar between co-FT and RETAIN-co-FT, while the benefit provided by RETAIN on OOD greatly increase with the generalist capabilities of the base model, and scales we pre-training data.
Figure 15: Detailed results on the benefits of RETAIN with increasing pretraining data.
A.4 Ablation on Learning Rate and Gradient Steps
Here, we ablate the learning rate and number of gradient steps we take in the task-FT policy in Fig.˜4 to study whether better hyperparameter choices can reduce or resolve overfitting [^1] Following the best practice from Black et al. [8], we always use a learning rate warmup period of $1000$ steps in LIBERO and a consine decay schedule over $30k$ steps to $1/10$ of the peak learning rate. The learning rate we report here is the peak learning rate. . In Fig.˜4, we use learning rate 2.5e52.5e-5. Here in Fig.˜16, we ablate four different learning rates: one greater than the original and two smaller. We evaluate the model at every $100$ gradient steps to also ablate on the number of gradient steps we take. With a larger learning rate, it’s clear that the overfitting issue is more severe, and the performance on all three kinds of evaluations (ID, OOD, Generalist) go down to near zero. With a smaller learning rate, the model still performs well in ID evaluations, and suffers less from forgetting generalist capabilities (measured from Generalist evaluations). This is to be expected because the finetuned model is closer to the pretrained model with a smaller learning rate. However, note that with a smaller learning rate, the OOD evaluation performance is also worse. We plot the OOD performance achieved using the original learning rate in Fig.˜4 as the dotted orange line in Fig.˜16, and the gap between the dotted and solid orange line shows the performance gap in OOD evaluation when we lower the learning rate, possibly due to underfitting. This experiment shows that tuning the learning rate and the number of update steps does not solve the problem of overfitting during finetuning. While lowering the learning rate can retain more generalist knowledge, it does not prevent the gradual loss of it. More importantly, the lower learning rate leads to worse OOD evaluation performance.
@@ A.4 Ablation on Learning Rate and Gradient Steps
A.5 Analysis of Finetuning Path in Parameter Space
@@
To start out, we check the colinearity of the vectors $\theta_{i+1}-\theta_{i}$ and $\theta_{i}-\theta_{i-1}$, where $\theta_{i}$ is the parameter of the finetuned checkpoint at gradient steps 100i100*i (i.e. we plot the difference vector every $100$ steps during finetuning). We measure the colinearity as the cosine similarity between the two difference vectors. Fig.˜17 shows this for the four different learning rates we ablated in Fig.˜16. Since no values are close to $1$, this shows that the changes in parameter space is highly non-linear. This is expected since the parameter trajectory of deep neural networks is often highly non-linear.
@@
Next, we attempt to more directly visualize the path/direction of the parameters during finetuning by projecting them down to 2D with Principal Component Analysis (PCA). Specifically, we again consider the difference of the weight vectors, $X_{i}=\theta_{i+1}-\theta_{i}$, at every $100$ steps during finetuning. In Fig.˜18, we plot the first two principal components of $X_{i}$ in blue, and label the points $i$. Indeed, we see that for all learning rates, the direction of the parameters is highly non-linear. With small learning rates, the direction oscillates a lot; with larger learning rates, the direction bends in a certain direction. In addition, we plot the two principal components of the parameter-merged model as well in orange. As expected, since the the merged model takes a linear path and the finetuned one does not, the two models end in in very different places in the parameter space. This shows that model merging achieves a different solution than any checkpoints on the finetuning path.
Finally, we analyze changes in all directions of the difference vectors $X_{i}$, instead of just the two principal components as shown in Fig.˜18. We take the difference vector matrix $Y=[X_{1};X_{2};...,X_{n}]$ and compute the singular values of $YY^{T}$ and Fig.˜19. If the parameter vectors lie in a linear path, then all difference vectors would point in the same direction and there should only be $1$ non-zero singular value. However, it’s clear from Fig.˜19 that most singular values are non-zero, showing that the path is non-linear in many dimensions. This generalizes the intuition from Fig.˜18 to more dimensions, and shows that model merging indeed achieves a different solution than the finetuning path.
@@
A.6 DROID Setup Details
@@ A.6.4 Evaluation Details
A.6.5 Evaluation Criteria: Whiteboard
@@
A.6.6 Evaluation Criteria: Plates
@@
A.6.7 Generalist Evaluations Details
@@
A.7 LIBERO Setup Details
@@
A.7.1 LIBERO Pretraining
In order to obtain a base-model to serve as the starting point for RETAIN in LIBERO, we pretrain $\pi_{0}$ on a mixture of LIBERO datasets. Specifically, we use 90 tasks from LIBERO-90, 9 tasks from LIBERO-object, 9 tasks from LIBERO-spatial, and 9 tasks from LIBERO-goal, for a total of 117 tasks in our pretraining dataset. For all LIBERO datasets, pretraining and finetuning, we utilize pre-processed RLDS versions gathered from here and here. These datasets consist of LIBERO demonstrations that have been preprocessed to upscale images, filter out transitions with idle actions, and remove failure trajectories. More details, such as hyperparameter choices, for the pretraining stage itself can be found in A.8. All subsequent fine-tuning and evaluation uses the normalization stats of this pretraining dataset, applying mean/standard-deviation normalization.
A.7.2 LIBERO Finetuning Datasets
As mentioned in 6.1, we finetune on 3 tasks from LIBERO-10: pot-on-stove, mugs-on-plates, and items-into-basket. For each dataset, we obtain pre-processed and filtered versions as described above. 20 highlights what these finetuning tasks, and thus also our ID evals, look like.